OpenStudy (anonymous):

the fastest a human has ever run is 27 miles per hours. How many miles per minute did the human run?

OpenStudy (anonymous):

0.45 miles per minute 27m/1hr 1hr = 60 min 27m/60min = xm/1min 27/60=0.45